40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 29 05:48:52 2011, kiwamu, Update, LSC, Y arm locked DSC_2978_ss.jpg
    Reply  Fri Apr 29 12:18:48 2011, kiwamu, Update, LSC, Y arm locked : details 
       Reply  Fri Apr 29 13:15:04 2011, kiwamu, Update, LSC, Y arm locked : details 
       Reply  Fri Apr 29 13:58:36 2011, rana, Update, LSC, Y arm locked : details 
Message ID: 4587     Entry time: Fri Apr 29 12:18:48 2011     In reply to: 4586     Reply to this: 4588   4589
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Y arm locked : details 

First of all, the conclusions / results from the exercise of the Y arm locking yesterday are:

   The position of the beam spots on both ETMY and ITMY are now not so bad ( ~ 5 mm off from the center).

   The input PZTs are coarsely aligned to the Y arm.

   Nevertheless IP_ANG is still too high to come out from the view port at the Y end station.

  After the alignments of PRM, SRM and Michelson, POP is still largely clipped.

 


(what I did)

  - Alignments of the Input PZTs

    First I tried letting the incident beam hit the center of ETMY by steering PZT1 and 2 as usual.
    Then I coarsely aligned the cavities to the incident beam and checked the beam flashing spots on ETMY and ITMY with the CCD monitors.
    When the spots were far from the center I went back to the alignment of PZT1 and 2 to get better beam positions. And repeated this work several times.

  - Adjustment of the demodulation phase  for the Y arm PDH.

    First I started looking at the digital signals and tried correcting the demodulation phase by the rotation matrix, but this didn't go fast because I had to do some DAQ settings, plotting and analysis.
    Instead looking at the digital signal, I observed the analog signals with an oscilloscope. I found the demodulation phase was something like 45 deg.
    Based on the analog measurement I rotated the digital matrix by 45 deg to get the I-signal maximized. Indeed this worked well. I obtained a beautiful PDH signal from the I-signal.
    Note that we are using 11 MHz mod/demod and eventually the signals come out from "REFL33" on the digital side.

  - Activation of oplev on ITMY

    Instead using ETMY I used ITMY for the length control because somehow I felt that ETMY coils were suspicious and they looked not so nice.
    One of the reason is that ETMY's coil actuation efficiencies looked low compared to the other test masses.
    For example a gain of 700 for SUSSIDE damping is needed on ETMY to get a reasonable Q. This is about 2 - 3 times larger gain than the other test masses.
    So I started using ITMY for the locking and activated the oplev to suppress unwanted excitation due to kicks from the control signal during the locking,
    The oplev has been misaligned, so I went to the ITMY optical bench and tweaked a steering mirror to let the He-Ne beam go into the QPD.
    I set the gains +2 for PITCH and -2 for YAW.

  - PDH locking

    The locking had been quite difficult even though the cavity alignment was quite good.
    It's because the beam on AS11_RFPD was almost falling off from the photo diode. This causes a big amplitude fluctuation in the PDH signal as the beam position moves.
    After aligning the beam by steering BS I got able to lock the Y arm. The PD whitening gains are all 0 dB and the feedback gain is -2, giving us a UGF of 250 Hz.

 

(Broken or likely broken stuff)

 * IP_ANG doesn't give a signal to the digital side.

 * ETMY coils look weak and 2 - 3 times weaker than the other test masses. (or OSEM readout gain maybe lower)

 * reload button on sitemap.adl doesn't work.

 * Farfalla, a lab laptop, seems out of network.

Quote from #4586

The Y arm has been locked with the IR beam. The purpose is to use the arm as an alignment reference for the input PZTs.

Detail will be posted later. Here is a picture of ITMY suspension. You can see there is a beam spot in the middle of the test mass.

 

ELOG V3.1.3-